# Robust Optimal Safe and Stability Guaranteeing Reinforcement Learning Control for Quadcopter

### v1

@@ V-A Reference Position Tracking
Figure 5: Control Input
Figure 6 demonstrates the controller performance in terms of quadcopter position, velocity, and its positional error. In the nominal controller-only scenario (Figure 6(a), (c), (e)), the quadcopter successfully stabilized at the reference point, but with slower error reduction. When combined with the NN controller (Figure 6(b), (d), (f)), the positional error decreased more rapidly, indicating improved transient performance. The NN controller effectively augmented the nominal controller by providing additional control inputs, leading to enhanced tracking accuracy and reduced response time, even under parameter variation.
@@
Figure 6: Results of reference point tracking
@@ V-B Reference Trajectory Tracking
To evaluate reference trajectory tracking performance, user-specified timed waypoints on an $S$-shaped trajectory was used as the initial specification, as listed in Table V. A path-planning algorithm of minimum snap method [18] was then used to map the timed waypoints of Table V to generate a smooth reference trajectory as a discrete time function for position, velocity, acceleration, and direction for trajectory tracking—See the red trajectory in Figure 7(a). During the control, the target position at each discrete sample time was adjusted to be the path planner’s computed position for the next sample instant, and our designed controller was employed for computing the new control action by simply changing the target position from that of the current sample instant to that of the next sample instant.
Figure 7 illustrates the quadcopter’s performance in tracking the $S$-shaped trajectory, where in Figure 7(a) the reference vs. actual trajectory are shown in red vs. black. The quadcopter successfully followed the reference trajectory while maintaining positional accuracy. Positional error remained low throughout the trajectory, validating the effectiveness of the NN controller in improving trajectory tracking precision even under parameter variations. The result demonstrates that the proposed control architecture is robust to environment uncertainties and capable of trajectory tracking.
TABLE V: Trajectory
@@
(a) Trajectory
(b) Position
(c) Velocity
(d) Positional Error
Figure 7: Results of trajectory tracking
@@ VI Conclusion
This paper presented certain improvements in the Robust Optimal Safe and Stability Guaranteed Training (ROSS-GT) method of [1] and its successful application for control of a quadcopter for reference trajectory tracking, in the presence of thrust and torque uncertainties representing the real-world disturbances.
The evaluation showed that it is feasible to design a robust, safe, and stability guaranteed controller for the quadcopter that also minimizes control cost and trajectory deviation, by combining the designed nominal cascaded PD controller with the designed NN controller. The addition of the NN controller further enhanced the transient/stead-state responses, demonstrating its capability to handle parameter variations effectively. The use of the robust stability-guaranteed approach ensured that the system remained asymptotically stable inspite of the parameter variations within the defined bounds.
Future work may focus on extending the method to more complex UAV models and high-fidelity simulation or real-world testing scenarios, further validating the applicability of stability-guaranteed NN controllers in practical environments.

### v2

@@ V Evaluation of Synthesized Controller
Figure 5: Reference position control input
@@
Figure 6: Results of reference position tracking
(a) Trajectory
(b) Position
(c) Velocity
(d) Positional Error
Figure 7: Results of reference trajectory tracking
V-A Reference position tracking
The reference position tracking capability of the quadcopter was evaluated under two scenarios: (1) using only the nominal controller and (2) using both the nominal and NN controllers. The quadcopter was initialized at $p_{init}=[0,0,0]^{T}$ with zero initial attitude, linear and angular velocities, while the reference linear position was set to $p_{d}=[1.4,0,0]^{T}$.
Control inputs from the nominal controller, the NN controller, and their combined values are shown in Figure 5. These inputs highlight the NN controller’s role in dynamically adjusting the system’s behavior to minimize position errors.
Figure 6 demonstrates the controller performance in terms of quadcopter position, velocity, and positional error. In the nominal controller-only scenario (Figure 6(a), (c), (e)), the quadcopter successfully stabilized at the reference point but with slower error reduction. When combined with the NN controller (Figure 6(b), (d), (f)), the positional error decreased much more rapidly, indicating improved transient performance. The NN controller effectively augmented the nominal controller by providing additional control inputs, leading to enhanced tracking accuracy and reduced response time, even under parameter variations.
@@ V-B Reference trajectory tracking
To evaluate reference trajectory tracking performance, user-specified timed waypoints on an $S$-shaped trajectory were used as the initial specification, as listed in Table V. A path-planning algorithm of the minimum snap method [26] was then used to map the timed waypoints of Table V to generate a smooth reference trajectory as a discrete time function for position, velocity, acceleration, and direction—See the red trajectory in Figure 7(a). During the control, the target position at each discrete sample time was set to be the path planner’s computed position for the next sample instant, and our designed controller was employed for computing the revised control action by simply changing the target position from that of the current sample instant to that of the next sample instant.
TABLE VI: Timed waypoints for a reference trajectory
@@
Figure 7 illustrates the quadcopter’s performance in tracking the $S$-shaped trajectory, where the reference vs. the actual trajectories are shown in red vs. black in Figure 7(a). The quadcopter successfully followed the reference trajectory by maintaining positional accuracy: Positional error remained low throughout the trajectory, validating the effectiveness of the NN controller in maintaining trajectory tracking precision even under parameter variations. The result demonstrates that the proposed control design is robust to environmental uncertainties and capable of trajectory tracking.
@@ VI Conclusion
This paper proposed novel improvements to the Robust Optimal Safe and Stability Guaranteed Control (ROSS-GC) method of [1], which enabled it for a successful first-time application for the control of a quadcopter for reference trajectory tracking in the presence of thrust and torque uncertainties, as is the case with the real-world scenarios. The developed controller with the improved design methodology showed the feasibility of a robust, optimal, safe, and stability-guaranteed controller synthesis for a quadcopter for the first time. The optimality stems from the minimization of trajectory deviation. The method involved combining a designed nominal cascaded PD controller with a NN controller—The former was designed to stabilize the nominal linearized system under no parametric changes, whereas the added NN control stabilized the given nonlinear system with parametric uncertainty. The addition of the NN controller demonstratively enhanced the transient/steady-state responses, showing its capability to handle parameter variations effectively, while the use of the robust stability-guaranteed approach certified the system’s asymptotic stability in spite of the parameter variations within the specified bounds. Future work can examine extending the method to more complex UAV models and validation through high-fidelity simulations as well as real-world testing, thereby further validating the general applicability of the proposed robust, optimal, safe, and stability-guaranteed NN controllers in practical environments.